OpenStudy (anonymous):

Solve -2log8(x + 1) = -8. Solve x = log94. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

@Abhishek619

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

\[-2\log_{8}(x+1)=-8\] multiply throughout by the (-1) \[2\log_{8} (x+1)=4\] then divide throughout by 2 \[\log_{8} (x+1)=4\] if you have reached this far you can then say

OpenStudy (jhannybean):

how did you get 4 on the RHS just by multiplying (-1)?

OpenStudy (primeralph):